OpenStudy (anonymous):

solve x+3 divided by 2=3x-5 divided by 10 i tried cross multiplying but i am not getting an answer that matches the multiple choices of either -10, 2,10 or -2??? plz help me

OpenStudy (campbell_st):

ok so you have \[\frac{x + 3}{2} = \frac{3x -5}{10}\] if you cross multiply you have \[10(x + 3) = 2(3x - 5)\] you need to distribute both sides \[10x + 30 = 6x - 10\] now solve for x

OpenStudy (anonymous):

yes i did all of that but the answer doesnt match 10, -10,2 nor -2 which are the multiple choice answers so which one is it

OpenStudy (campbell_st):

your arithmetic is incorrect as the answer is in your choices start by subtracting 6x from both sides, what do you get...?

OpenStudy (anonymous):

yes you are right thnak you answer

OpenStudy (anonymous):

it is -10